# Low-temperature transport in high-conductivity correlated metals: a density-functional plus dynamical mean-field study of cubic perovskites

### v1

@@ Abstract
While methods based on density-functional perturbation theory have dramatically improved our understanding of electron-phonon contributions to transport in materials, methods for accurately capturing electron-electron scattering relevant to low temperatures have seen significantly less development. The case of high-conductivity, moderately correlated materials characterized by low scattering rates is particularly challenging, since exquisite numerical precision of the low-energy electronic structure is required. Recent methodological advancements to density-functional theory combined with dynamical mean-field theory (DFT+DMFT), including adaptive Brillouin-zone integration and numerically precise self-energies, enable a rigorous investigation of electron-electron scattering in such materials. In particular, these tools may be leveraged to perform a robust scattering-rate analysis on both real- and imaginary-frequency axes. Applying this methodology to a subset of ABO3 perovskite oxides – SrVO3, SrMoO3, PbMoO3, and SrRuO3 – we demonstrate its ability to qualitatively and quantitatively describe electron-electron contributions to the temperature-dependent direct-current resistivity. This combination of numerical techniques offers fundamental insight into the role of electronic correlations in transport phenomena and provides a predictive tool for identifying materials with potential for technological applications.
@@ I Introduction
We deploy our computational framework on a representative subset of ABO3 perovskite oxides: SrVO3, SrMoO3, PbMoO3, and SrRuO3. These materials were chosen since they have a similar atomic and electronic structure but distinct transport properties observed experimentally. SrVO3 and SrMoO3 are moderately correlated high-conductivity metals, with $d^{1}$ and $d^{2}$ occupancies of their $t_{2g}$ orbitals, respectively. In fact, SrMoO3 has the lowest reported room-temperature (RT) resistivity of any perovskite oxide [32]. As we demonstrate in our companion paper [33], resistivity measurements on both SrVO3 and SrMoO3 are consistent with FL $\rho(T)\propto T^{2}$ at low $T$, as well as approximately $T^{2}$ el-ph contributions around RT [7, 34]. PbMoO3 has the same formal valence as SrMoO3 but a measured $\rho$ three orders of magnitude higher and an unusual sublinear $T$ dependence [35]. Finally, SrRuO3 ($d^{4}$) is strongly correlated and will be used to contrast with the previous materials. It exhibits non-FL transport down to 10similar-toabsent10$\sim$ 10∼ 1030303030 K [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Experimentally, the crystal structure of these oxides is either cubic at all $T$ (SrVO3) or exhibits a cubic to orthorhombic transition as $T$ is lowered. In order to make direct comparisons between the different materials, we employ the cubic ($Pm\overline{3}m$) crystal structure for all of them.
The focus of this work is on the el-el contributions to $\rho(T)$. For moderately correlated metals, one expects that el-ph scattering, which is not considered here, dominates at all $T$ relevant for practical applications (i.e., around RT). However, achieving a quantitative understanding of el-el scattering remains crucial—not only for fundamental reasons but also, as demonstrated in Ref. 33, to clarify transport behavior by ruling out el-el scattering as the origin of the observed $\rho$ versus $T$ scaling at RT.
@@
The remainder of this paper is organized as follows. In Section II, we present the Kubo transport formalism and subsequently derive an expression for $\rho$ using the FL form of the electronic self-energy. In Section III, we describe various components of our computational methodology to compute $\rho$ within DFT+DMFT. In Section IV, we present our computed $\rho(T)$ for each material. Section V contains a summary of the paper and discusses the implications and outlook.
@@ II Transport formalism
As is common in DFT+DMFT calculations of the DC resistivity, we start from the Kubo formula for the conductivity. In this work, we focus on cubic oxides involving the three degenerate $t_{2g}$ $d$ orbitals. Then, within the DMFT approximation of a local self-energy, the self-energy $\Sigma_{\omega}$ is a scalar (proportional to the unit matrix). It follows that, consistent with the momentum independence of the self-energy and (irreducible) vertex, vertex corrections are absent in DMFT [47, 48], so that
@@
In the coherent regime, where the spectral function is sharp compared to temperature, intraband contributions dominate the DC conductivity, so that
@@
Using $[\Sigma_{\omega}]_{\nu\nu^{\prime}}=\delta_{\nu\nu^{\prime}}\Sigma_{\omega}$, we can express the spectral function as
@@
where $\mu$ is the chemical potential and $\epsilon_{\bm{k}\nu}$ are the DFT eigenvalues. Since $[A_{\bm{k}\omega}]_{\nu\nu}$ depends on $\bm{k}$ only via $\epsilon_{\bm{k}\nu}$, we can write the conductivity ($\sigma=\sigma^{\alpha\alpha}$, $\alpha\in\{x,y,z\}$) as
@@
In the last step, we expanded $\Phi_{\omega+E^{\mathrm{F}}_{\omega}+\frac{1}{2}\Gamma_{\omega}y}$ in $y$, using that odd orders vanish under the integral.
@@
The remaining calculation involves only the scattering rate, which we take in FL form as
@@
Thus, we have two options for calculating the FL $\rho(T)$ at low $T$, the full Kubo formula in Eq. (1) or the simplified version of Eq. (12). While more general, Eq. (1) involves the momentum-resolved real-frequency spectral function at low frequencies, which can be challenging to obtain accurately for materials with low scattering rates. This issue is partially addressed by using adaptive momentum-integration techniques (Sec. III.2), but also stems from problems with analytic continuation of the self-energy at low $\omega$ (Sec. III.6). Alternatively, we will show in Sec. III.6 that the constant $C$ in Eq. (12) can be obtained from higher-frequency ranges of the FL self-energy utilizing the known FL scaling [Eq. (10)]. This makes Eq. (12) a robust choice for obtaining $\rho$ in high-conductivity materials at low $T$. We compare both options in Sec. IV.
@@ III.1 Interpolating DFT
These transition-metal $t_{2g}$ states will serve as the active correlated subspace for our DMFT calculations. We downfold onto these bands by constructing a local atomic-like orbitals using maximally localized Wannier functions (MLWF) [55] as implemented in Wannier90 [56]. The band structure obtained from our Wannier functions is compared to the Kohn–Sham (KS) DFT band structure in Fig. 1(b). The agreement demonstrates that our Wannier functions are a faithful representation of the DFT bands near the Fermi level. In order to obtain a consistent description for all four materials, we have disentangled the $t_{2g}$ bands from any other hybridizing DFT bands. Table 1 provides the window parameters used for each material. The width of the “frozen windows,” where exact energy agreement is enforced, are indicated in light purple in Fig. 1(b).
Table 1: Energy windows for the calculation of Wannier functions. All energies are given in eV and normalized to the Fermi level for each material.
@@
SrVO3 $0.907$ $-1.192$ $3.007$ $-1.292$
SrMoO3 $1.768$ $-2.531$
PbMoO3 $-0.094$ $-1.994$ $3.705$ $-3.494$
SrRuO3 $0.028$ $-2.171$ $0.828$ $-3.171$
Figure 2: Convergence of SrMoO3 DFT calculated properties at the Fermi level $\epsilon_{\mathrm{F}}$. (a) DOS and transport function for various broadening $\eta$, obtained with the iterated adaptive integrator (IAI). Relative differences are taken from the values $V_{\mathrm{uc}}D(\epsilon_{\mathrm{F}})=1.886$/eV, $\Phi(\epsilon_{\mathrm{F}})=3.930\,$eV/($\Omega\,$cm). (b) DOS at each value of $\eta$ computed with a fixed number of $k$ points, $N_{k}$, per dimension in the full BZ. Differences are taken from the IAI result at the respective value of $\eta$.
III.2 Brillouin-zone integration
@@
where $A^{\eta}_{\bm{k}\omega}$ is the broadened non-interacting spectral function, which can be evaluated in the orbital basis as
@@
$[A^{\eta}_{\bm{k}\omega}]_{\nu\nu}[A^{\eta}_{\bm{k}\omega}]_{\nu^{\prime}\nu^{\prime}}\approx\frac{\eta}{\pi}\frac{\delta(\omega-\epsilon_{\bm{k}\nu})+\delta(\omega-\epsilon_{\bm{k}\nu^{\prime}})}{(\epsilon_{\bm{k}\nu}-\epsilon_{\bm{k}\nu^{\prime}})^{2}+(2\eta)^{2}},$ (15)
implying $[A^{\eta}_{\bm{k}\omega}]_{\nu\nu}^{2}\approx\frac{1}{2\pi\eta}\delta(\omega-\epsilon_{\bm{k}\nu})$. Interband contributions, absent in Eq. (6), are thus suppressed by $\mathit{O}(\eta^{2})$, and the desired result follows in the limit $\eta\to 0$.
Figure 2(a) demonstrates the convergence of the DOS and transport function at the Fermi level with respect to $\eta$, compared with the result at $\eta=1$ meV rounded to three digits (as used in Ref. 33). We observe that convergence to below $10^{-3}$ ($10^{-4}$) is achieved at $\eta=8$ meV ($\eta=2$ meV) for the DOS (transport function). In Fig. 2(b), we plot the error of the uniform integration method implemented in AutoBZ.jl versus the number $N_{k}$ of quadrature points per dimension in the full BZ. Convergence to below $10^{-2}$ requires about $N_{k}=200$ at $\eta=8$ meV and about $N_{k}=800$ at $\eta=1$ meV. Note that this requires $N_{k}^{3}$ grid points to resolve the BZ integral. Finally, we plot the calculated $t_{2g}$ DOS and transport function for our oxide materials using IAI with $\eta=1$ meV in Fig. 3.
@@
The next step is to determine the interaction parameters that define the Hubbard–Kanamori Hamiltonian of the correlated $t_{2g}$ impurity problem identified above. Ideally, these parameters are derived from first principles, avoiding any empirical input. The constrained random phase approximation (cRPA) [72] is the standard approach to estimate the static interaction parameters from DFT. For our test case oxides, we compute cRPA interaction parameters utilizing the implementation in the VASP code [73] by averaging the full four-index interaction tensor to parameterize the Hubbard–Kanamori Hamiltonian [74]. By the cubic symmetry within the $t_{2g}$ subspace, these parameters reduce to a single Hubbard $U$ and Hund’s coupling $J$ parameter, which are summarized for each material in Table 2.
@@
Despite its successes in many cases, cRPA can sometimes fail to capture key experimental observations; this is the case for SrVO3 and SrRuO3. For SrVO3, experiments report a quasiparticle renormalization factor of $Z\sim 0.5$ [76]. In DMFT, this quantity can be estimated from the self-energy as $Z^{-1}=1-\partial_{\omega}\mathrm{Re}\Sigma(\omega,T)|_{\omega\rightarrow 0}$. Using the cRPA-derived paramaters in DMFT, we obtain $Z\sim 0.6$, while increasing $(U,J)$ to (4.5, 0.65) yields $Z=0.5$, in agreement with experiment [77].
For SrRuO3, the cRPA-predicted interaction parameters fail to capture the correct behavior with respect to the ferromagnetic transition. This transition is especially sensitive to $J$. For the cRPA value of 0.26 eV, we do not observe a ferromagnetic transition down to the lowest temperature studied, while increasing the value to 0.40 eV results in a transition around 290 K. Experimentally, the value is around 150-170 K [36, 78, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], this overestimation is expected since DMFT only includes local magnetic fluctuations and neglects their spatial dependence [1]—see, e.g., Ref. 79 for DMFT estimates of the Curie temperature of transition metals. We note that the interaction parameters are significantly constrained by the fact that they should be roughly the same for SrRuO3 and CaRuO3, and produce a ferromagnetic (paramagnetic) ground state at low temperatures for SrRuO3 (CaRuO3) [80]. This is indeed satisfied for our choice of $U=2.35$ eV and $J=0.4$ eV, see Ref. 80.
@@ III.4 Electron-electron self-energies from DMFT
We solve the quantum impurity model for the $t_{2g}$ correlated subspace of each material to obtain the self-energy from el-el interactions within DMFT. The DC resistivity depends sensitively on the low-frequency behavior of $\Sigma(\omega)$ [see Eq. (1)] requiring accurate real-frequency data. Commonly used impurity solvers are based on continuous-time quantum Monte Carlo (CT-QMC), which operates exclusively in imaginary time. To obtain real-frequency data for Eq. (1), one must analytically continue from the imaginary axis to the real axis, which is an ill-conditioned mathematical problem [81] and therefore prone to errors and artifacts.
@@
As shown in Fig. 4, we obtain a “handshake” agreement between the two impurity solvers on both the imaginary- (top row) and real-frequency axes (bottom row) for all materials. SrVO3, SrMoO3, and PbMoO3 all show the characteristic FL scaling, i.e., $-\mathrm{Im}\Sigma(\omega)\propto\omega^{2}+\pi^{2}T^{2}$ at low energies, while SrRuO3 clearly exhibits non-FL behavior at this temperature ($T=116$ K). For QMC, this agreement gives us confidence in the convergence of the imaginary-frequency self-energies as well as the AC (at least at relatively low frequencies). For NRG, this system (three orbitals, full Hubbard–Kanamori Hamiltonian with spin-flip and pair-hopping terms) is a considerable challenge, and resolving the FL behavior in $\mathrm{Im}\Sigma$ has only recently become possible [112]. Thus, the agreement in Fig. 4 with the QMC data also gives confidence that the NRG self-energies are accurate.

### v2

@@ Abstract
While methods based on density-functional perturbation theory have dramatically improved our understanding of electron-phonon contributions to transport in materials, methods for accurately capturing electron-electron scattering relevant to low temperatures have seen significantly less development. The case of high-conductivity, moderately correlated materials characterized by low scattering rates is particularly challenging, since exquisite numerical precision of the low-energy electronic structure is required. Recent methodological advancements to density-functional theory combined with dynamical mean-field theory (DFT+DMFT), including adaptive Brillouin-zone integration and numerically precise self-energies, enable a rigorous investigation of electron-electron scattering in such materials. In particular, these tools may be leveraged to perform a robust scattering-rate analysis on both real- and imaginary-frequency axes. Applying this methodology to a subset of ABO3 perovskite oxides – SrVO3, SrMoO3, PbMoO3, and SrRuO3 – we demonstrate its ability to obtain quantitative convergence of the local electron-electron contributions to the temperature-dependent direct-current resistivity. This combination of numerical techniques offers fundamental insight into the role of electronic correlations in transport phenomena and provides a predictive tool for identifying materials with potential for technological applications.
@@ I Introduction
We deploy our computational framework on a representative subset of ABO3 perovskite oxides: SrVO3, SrMoO3, PbMoO3, and SrRuO3. These materials were chosen because they have a similar atomic and electronic structure but distinct transport properties observed experimentally. SrVO3 and SrMoO3 are moderately correlated high-conductivity metals, with $d^{1}$ and $d^{2}$ occupancies of their $t_{2g}$ orbitals, respectively. In fact, SrMoO3 has the lowest reported room-temperature (RT) resistivity of any perovskite oxide [32]. As we demonstrate in our companion paper [33], resistivity measurements on both SrVO3 and SrMoO3 are consistent with FL $\rho(T)\propto T^{2}$ at low $T$, as well as approximately $T^{2}$ el-ph contributions around RT [7, 34]. PbMoO3 has the same formal valence as SrMoO3 but a measured $\rho$ three orders of magnitude higher and an unusual sublinear $T$ dependence [35]. Finally, SrRuO3 ($d^{4}$) is strongly correlated and will be used to contrast with the previous materials. It exhibits non-FL transport down to $\sim 10$$30$ K [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Experimentally, the crystal structure of these oxides is either cubic at all $T$ (SrVO3) or exhibits a cubic to orthorhombic transition as $T$ is lowered. To make direct comparisons between the different materials, we employ the cubic ($Pm\overline{3}m$) crystal structure for all of them.
The focus of this work is on the local el-el contributions to $\rho(T)$. For moderately correlated metals, one expects that el-ph scattering, which is not considered here, dominates at all $T$ relevant for practical applications (i.e., around RT). However, achieving a quantitative understanding of el-el scattering remains crucial—not only for fundamental reasons but also, as demonstrated in Ref. 33, to clarify transport behavior by ruling out el-el scattering as the origin of the observed $\rho$ versus $T$ scaling at RT.
@@
The remainder of this paper is organized as follows. In Sec. II, we present the Kubo transport formalism and subsequently derive an expression for $\rho$ using the FL form of the electronic self-energy. In Sec. III, we describe various components of our computational methodology to compute $\rho$ within DFT+DMFT. In Sec. IV, we present our computed $\rho(T)$ for each material. Section V contains a summary of the paper and discusses the implications and outlook.
@@ II Transport formalism
As is common in DFT+DMFT calculations of the dc resistivity, we start from the Kubo formula for the conductivity. In this work, we focus on cubic oxides involving the three degenerate $t_{2g}$ $d$ orbitals. Then, within the DMFT approximation of a local self-energy, the self-energy $\Sigma_{\omega}$ is a scalar (proportional to the unit matrix). It follows that, consistent with the momentum independence of the self-energy and (irreducible) vertex, vertex corrections are absent in DMFT [47, 48], so
@@
In the coherent regime, where the spectral function is sharp compared to temperature, intraband contributions dominate the dc conductivity, so
@@
Using $\Sigma_{\omega}]_{\nu\nu^{\prime}}=\delta_{\nu\nu^{\prime}}\Sigma_{\omega}$, we can express the spectral function as
@@
where $\mu$ is the chemical potential and $\epsilon_{\bm{k}\nu}$ are the DFT eigenvalues. Since $A_{\bm{k}\omega}]_{\nu\nu}$ depends on $\bm{k}$ only via $\epsilon_{\bm{k}\nu}$, we can write the conductivity ($\sigma=\sigma^{\alpha\alpha}$, $\alpha\in\{x,y,z\}$) as
@@
$\Phi_{\omega+E^{\mathrm{F}}_{\omega}+\frac{1}{2}\Gamma_{\omega}y}$ in $y$, using that odd orders vanish under the integral.
@@
Note that higher orders in the expansion of $\Phi$ lead to higher orders beyond $T^{2}$ in the expansion of $\rho$ and are thus irrelevant for the present derivation. The remaining calculation involves only the scattering rate, which we take in FL form as
@@
Thus, we have two options for calculating the FL $\rho(T)$ at low $T$, the full Kubo formula in Eq. (1) or the simplified version of Eq. (12). While more general, Eq. (1) involves the momentum-resolved real-frequency spectral function at low frequencies, which can be challenging to obtain accurately for materials with low scattering rates. This issue is partially addressed by using adaptive momentum-integration techniques (Sec. III.2), but also stems from problems with analytic continuation of the self-energy at low $\omega$ (Sec. III.6). Alternatively, we will show in Sec. III.6 that the constant $C$ in Eq. (12) can be obtained from higher-frequency ranges of the FL self-energy utilizing the known FL scaling [Eq. (10)]. This makes Eq. (12) a robust choice for obtaining $\rho$ in high-conductivity materials at low $T$. We compare both options in Sec. IV (see Fig. 8).
@@ III.1 Interpolating DFT
These transition-metal $t_{2g}$ states will serve as the active correlated subspace for our DMFT calculations. We downfold onto these bands by constructing a local atomic-like orbitals using maximally localized Wannier functions (MLWF) [55] as implemented in Wannier90 [56]. The band structure obtained from our Wannier functions is compared to the Kohn–Sham DFT band structure in Fig. 1(b). The agreement demonstrates that our Wannier functions are a faithful representation of the DFT bands near the Fermi level. To obtain a consistent description for all four materials, we have disentangled the $t_{2g}$ bands from any other hybridizing DFT bands. Table 1 provides the window parameters used for each material. The width of the “frozen windows,” where exact energy agreement is enforced, are indicated in light purple in Fig. 1(b).
Table 1: Energy windows for the calculation of Wannier functions. All energies are given in eV and relative to the Fermi level for each material.
@@
SrVO3 $0.907$ 1.192-1.192 $3.007$ 1.292-1.292
SrMoO3 $1.768$ 2.531-2.531
PbMoO3 0.094-0.094 1.994-1.994 $3.705$ 3.494-3.494
SrRuO3 $0.028$ 2.171-2.171 $0.828$ 3.171-3.171
@@
where $A^{\eta}_{\bm{k}\omega}$ is the broadened noninteracting spectral function, which can be evaluated in the orbital basis as
@@
$A^{\eta}_{\bm{k}\omega}]_{\nu\nu}[A^{\eta}_{\bm{k}\omega}]_{\nu^{\prime}\nu^{\prime}}\approx\frac{\eta}{\pi}\frac{\delta(\omega-\epsilon_{\bm{k}\nu})+\delta(\omega-\epsilon_{\bm{k}\nu^{\prime}})}{(\epsilon_{\bm{k}\nu}-\epsilon_{\bm{k}\nu^{\prime}})^{2}+(2\eta)^{2}},$ (15)
implying $A^{\eta}_{\bm{k}\omega}]_{\nu\nu}^{2}\approx\frac{1}{2\pi\eta}\delta(\omega-\epsilon_{\bm{k}\nu})$. Interband contributions, absent in Eq. (6), are thus suppressed by $\mathit{O}(\eta^{2})$, and the desired result follows in the limit $\eta\to 0$.
Figure 2: Convergence of SrMoO3 DFT calculated properties at the Fermi level $\epsilon_{\mathrm{F}}$. (a) DOS and transport function for various broadening $\eta$, obtained with the iterated adaptive integrator (IAI). Relative differences are taken from the values $V_{\mathrm{uc}}D(\epsilon_{\mathrm{F}})=1.886$/eV, $\Phi(\epsilon_{\mathrm{F}})=3.930\,$eV/($\Omega\,$cm). (b) DOS (solid lines) and transport function (dotted lines) at each value of $\eta$ computed with a fixed number of $k$ points, $N_{k}$, per dimension in the full BZ. Differences are taken from the IAI result at the respective value of $\eta$.
III.2 Brillouin-zone integration
Figure 2(a) demonstrates the convergence of the DOS and transport function at the Fermi level with respect to $\eta$, compared with the result at $\eta=1$ meV rounded to three digits (as used in Ref. 33). We observe that convergence to below $10^{-3}$ ($10^{-4}$) is achieved at $\eta=8$ meV ($\eta=2$ meV) for the DOS (transport function). In Fig. 2(b), we plot the error of the uniform integration method for the DOS, implemented in AutoBZ.jl, versus the number $N_{k}$ of quadrature points per dimension in the full BZ. Convergence to below $10^{-2}$ requires about $N_{k}=200$ ($N_{k}=300$) at $\eta=8$ meV and about $N_{k}=700$ ($N_{k}=1000$) at $\eta=1$ meV for the DOS (transport function). Note that this requires $N_{k}^{3}$ grid points to resolve the BZ integral. Finally, we plot the calculated $t_{2g}$ DOS and transport function for our oxide materials using IAI with $\eta=1$ meV in Fig. 3.
@@ III.2 Brillouin-zone integration
The next step is to determine the interaction parameters that define the Hubbard–Kanamori Hamiltonian of the correlated $t_{2g}$ impurity problem identified above. Ideally, these parameters are derived from first principles, avoiding any empirical input. The constrained random phase approximation (cRPA) [72] is the standard approach to estimate the static interaction parameters from DFT. For our test case oxides, we compute cRPA interaction parameters utilizing the implementation in the VASP code [73] by averaging the full four-index interaction tensor to parametrize the Hubbard–Kanamori Hamiltonian [74]. By the cubic symmetry within the $t_{2g}$ subspace, these parameters reduce to a single Hubbard $U$ and Hund’s coupling $J$ parameter, which are summarized for each material in Table 2.
@@
Despite its successes in many cases, cRPA can sometimes fail to capture key experimental observations; this is the case for SrVO3 and SrRuO3. For SrVO3, experiments report a quasiparticle renormalization factor of $Z\sim 0.5$ [76]. In DMFT, this quantity can be estimated from the self-energy as $Z^{-1}=1-\partial_{\omega}\mathrm{Re}\Sigma(\omega,T)|_{\omega\rightarrow 0}$. Using the cRPA-derived paramaters in DMFT, we obtain $Z\sim 0.6$, while increasing $(U,J)$ to (4.5, 0.65) yields $Z=0.5$, in agreement with experiment [77]. This artificial increase of the static Hubbard interaction to correctly capture $Z$ has been discussed previously [77, 78]. It can be circumvented by including dynamical interactions in combination with non-local correlation effects [79, 80, 78].
Fol SrRuO3, the cRPA-predicted interaction parameters fail to capture the correct behavior with respect to the ferromagnetic transition. This transition is especially sensitive to $J$. For the cRPA value of 0.26 eV, we do not observe a ferromagnetic transition down to the lowest temperature studied, while increasing the value to 0.40 eV results in a transition around 290 K. Experimentally, the value is around 150-170 K [36, 81, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], this overestimation is expected since DMFT only includes local magnetic fluctuations and neglects their spatial dependence [1]—see, e.g., Ref. 82 for DMFT estimates of the Curie temperature of transition metals. We note that the interaction parameters are significantly constrained by the fact that they should be roughly the same for SrRuO3 and CaRuO3, and produce a ferromagnetic (paramagnetic) ground state at low temperatures for SrRuO3 (CaRuO3) [83]. This is indeed satisfied for our choice of $U=2.35$ eV and $J=0.4$ eV, see Ref. 83.
@@ III.4 Electron-electron self-energies from DMFT
We solve the quantum impurity model for the $t_{2g}$ correlated subspace of each material to obtain the local self-energy from el-el interactions within DMFT. The dc resistivity depends sensitively on the low-frequency behavior of $\Sigma(\omega)$ [see Eq. (1)] requiring accurate real-frequency data. Commonly used impurity solvers are based on continuous-time quantum Monte Carlo (CT-QMC), which operates exclusively in imaginary time. To obtain real-frequency data for Eq. (1), one must analytically continue from the imaginary axis to the real axis, which is an ill-conditioned mathematical problem [84] and therefore prone to errors and artifacts.
@@
As shown in Fig. 4, we obtain a “handshake” agreement between the two impurity solvers on both the imaginary- (top row) and real-frequency axes (bottom row) for all materials. SrVO3, SrMoO3, and PbMoO3 all show the characteristic FL scaling, i.e., $\mathrm{Im}\Sigma(\omega)\propto\omega^{2}+\pi^{2}T^{2}$ at low energies, while SrRuO3 clearly exhibits non-FL behavior at this temperature ($T=116$ K). For QMC, this agreement gives us confidence in the convergence of the imaginary-frequency self-energies as well as the AC (at least at relatively low frequencies). For NRG, this system (three orbitals, full Hubbard–Kanamori Hamiltonian with spin-flip and pair-hopping terms) is a considerable challenge, and resolving the FL behavior in $\mathrm{Im}\Sigma$ has only recently become possible [115]. Thus, the agreement in Fig. 4 with the QMC data also gives confidence that the NRG self-energies are accurate.